You are analyzing the returns of an investment fund over the last 2 years. Returns and inflation during that time are listed below in the table. What was the real return during year 2? Answer in percentage points to the nearest 4th decimal.

Return Inflation

Year 1 7.20% 3.90%

Year 2 2.10%. 4.20%
Answer #1

According to the Fisher Equation,

(1 + i) = (1 + r) × (1 + π)

where, i = nominal interest rate

r = real interest rate

π = inflation rate

Hence, r = (1+ i) / (1 + π) - 1

Year 1

i = 0.0720
π = 0.0390
=> r = (1 + 0.0720)/(1 + 0.0390) - 1 = 0.0318 or 3.18%

Year 1

i = 0.0210
π = 0.0420
=> r = (1 + 0.0420)/(1 + 0.0210) - 1 = 0.0206 or 2.06%